What is the maximum Truck Bed Height you can successfully load a 5 W piano on by 2 friends?


1
Expert's answer
2021-05-30T13:26:35-0400

Assume the two friends can perform a certain amount of work, call it "E". Then the height is defined by


"h=\frac{E}{W},"

where "W" - weight. So, if the piano weigh 5 times more, the height is


"h_1=\frac{E}{5W}=\frac h5."

The maximum Truck Bed Height is 5 times lower.